\section*{References}
\begin{thebibliography}{9}
%\cite{}
\bibitem{6d} 
  L.~D.~Roper,
  %``Evidence for a P-11 Pion-Nucleon Resonance at 556 MeV,''
  Phys.\ Rev.\ Lett.\  {\bf 12}, 340 (1964);
  %%CITATION = PRLTA,12,340;%%
  %91 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  G.~T.~Hoff,
  %``Possible resonance in 829 MeV Lambda K^0 production,''
  Phys.\ Rev.\ Lett.\  {\bf 12}, 652 (1964);
  %%CITATION = PRLTA,12,652;%%
  %9 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  J.~H.~Christenson, J.~W.~Cronin, V.~L.~Fitch and R.~Turlay,
  %``Evidence for the 2 pi Decay of the k(2)0 Meson,''
  Phys.\ Rev.\ Lett.\  {\bf 13}, 138 (1964).
  %%CITATION = PRLTA,13,138;%%
  %2285 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  H.~R.~Crouch, R.~Hargraves, B.~Kendall, R.~E.~Lanou, A.~M.~Shapiro, M.~Widgoff, G.~E.~Fischer and A.~E.~Brenner {\it et al.},
%  %``N-33* (1238) and rho0 Production by High-Energy Photons,''
%  Phys.\ Rev.\ Lett.\  {\bf 13}, 640 (1964);
% %%CITATION = PRLTA,13,640;%%
% %13 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  H.~C.~DeStaebler, E.~F.~Erickson, A.~C.~Hearn and C.~Schaerf,
  %``Photoproduction of $\pi^0$ from Hydrogen near the Second Pion-Nucleon Resonance,''
  Phys.\ Rev.\  {\bf 140}, B336 (1965);
  %%CITATION = PHRVA,140,B336;%%
  %24 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  G.~T.~Hoff,
%  %``Analysis and model of low-energy Lambda K^0 production,''
%  Phys.\ Rev.\  {\bf 139}, B671 (1965);
%  %%CITATION = PHRVA,139,B671;%%
% %5 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  B.~H.~Bransden, P.~J.~O'Donnell and R.~G.~Moorhouse,
  %``Pion-Nucleon Scattering Amplitudes in the Range 300-700 MeV,''
  Phys.\ Rev.\  {\bf 139}, B1566 (1965);
  %%CITATION = PHRVA,139,B1566;%%
  %34 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
 % J.~E.~Rush and W.~G.~Holladay,
 % %``Application of pi - N scattering results to Lambda K^0 production,''
 % Phys.\ Rev.\  {\bf 148}, 1444 (1966);
 % %%CITATION = PHRVA,148,1444;%%
 % %4 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  L.~Bertanza, A.~Bigi, R.~Carrara and R.~Casali,
  %``pi^- - p interactions at 775 MeV,''
  Nuovo Cim.\ A {\bf 44}, 712 (1966);
  %%CITATION = NUCIA,A44,712;%%
  %10 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  E.~W.~Anderson, E.~J.~Bleser, G.~B.~Collins, T.~Fujii, J.~Menes, F.~Turkot, R.~A.~Carrigan and R.~M.~Edelstein {\it et al.},
  %``Experimental Study of p+p --> p+N* at Incident Energies of 6-30 BeV,''
  Phys.\ Rev.\ Lett.\  {\bf 16}, 855 (1966);
  %%CITATION = PRLTA,16,855;%%
  %122 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  F.~F.~Liu and S.~Vitale,
%  %``Positive-Pion Production Asymmetry with Polarized Bremsstrahlung Near Second Resonance,''
%  Phys.\ Rev.\  {\bf 144}, 1093 (1966);
%  %%CITATION = PHRVA,144,1093;%%
% %11 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  K.~J.~Foley, R.~S.~Jones, S.~J.~Lindenbaum, W.~A.~Love, S.~Ozaki, E.~D.~Platner, C.~A.~Quarles and E.~H.~Willen,
  %``N* Production in pi + /- -p and p-p Interactions,''
  Phys.\ Rev.\ Lett.\  {\bf 19}, 397 (1967);
  %%CITATION = PRLTA,19,397;%%
  %51 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  T.~H.~Tan, M.~L.~Perl, F.~Martin, W.~Chinowsky, R.~Kinsey, S.~Klein, P.~Schmidt and M.~Mandelkern {\it et al.},
%  %``An Investigation Of The 1.4-gev/c**2 Nucleon Isobar In Proton Proton Interactions,''
%  Phys.\ Lett.\ B {\bf 28}, 195 (1969);
%  %%CITATION = PHLTA,B28,195;%%
% %4 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  C.~M.~Ankenbrandt, A.~R.~Clark, B.~Cork, T.~Elioff, L.~T.~Kerth and W.~A.~Wenzel,
  %``Nucleon Isobar Production in Proton-Proton Collisions between 3 and 7 GeVc,''
  Phys.\ Rev.\  {\bf 170}, 1223 (1968);
  %%CITATION = PHRVA,170,1223;%%
  %66 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  J.~W.~Lamsa, N.~M.~Cason, N.~N.~Biswas, I.~Derado, T.~H.~Groves, V.~P.~Kenney, J.~A.~Poirier and W.~D.~Shephard,
  %``Resonances Produced in the Reaction pi-p --> pi-pi-pi+p at 8 GeV/c,''
  Phys.\ Rev.\  {\bf 166}, 1395 (1968);
  %%CITATION = PHRVA,166,1395;%%
  %14 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  R.~J.~Miller, S.~Lichtman and R.~B.~Willmann,
  %``Four-charged-particle final states from the interaction of 2.7 gev/c pi-plus on deuterium,''
  Phys.\ Rev.\  {\bf 178}, 2061 (1969);
  %%CITATION = PHRVA,178,2061;%%
  %27 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  S.~P.~Almeida, J.~G.~Rushbrooke, J.~H.~Scharenguivel, M.~Behrens, V.~Blobel, I.~Borecka and H.~C.~Dehne {\it et al.},
  %``P-p interactions at 10 gev/c,''
  Phys.\ Rev.\  {\bf 174}, 1638 (1968);
  %%CITATION = PHRVA,174,1638;%%
  %54 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  T.~Ferbel, R.~Holmes and S.~Stone,
%  %``Production of low-mass p + pi-plus + pi-minus and anti-p + pi-plus + pi-minus states in the reaction anti-p + p ---> anti-p + p + pi-plus + pi-minus at 3-4 gev/c,''
%  Phys.\ Rev.\ Lett.\  {\bf 22}, 1141 (1969);
%  %%CITATION = PRLTA,22,1141;%%
% %4 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  T.~C.~Bacon, F.~Bomse, T.~B.~Borak, T.~B.~Cochran, W.~J.~Fickinger, E.~R.~Goza, H.~W.~K.~Hopkins and E.~O.~Salant,
  %``Evidence for production of an n*(1400) and anti-n*(1400) in the reaction anti-p + p(n) ---> anti-p + p + pi-plus + pi-minus(n) at 2.8 gev/c,''
  Phys.\ Rev.\ Lett.\  {\bf 22}, 43 (1969).
  %%CITATION = PRLTA,22,43;%%
  %9 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%%\bibitem{} 
%  A.~Shapira, O.~Benary, Y.~Eisenberg, E.~E.~Ronat, D.~Yaffe and G.~Yekutieli,
%  %``Production of the n*(1470) in p + d ---> p(s) + p + p + pi-minus at 7.0 gev/c,''
%  Phys.\ Rev.\ Lett.\  {\bf 21}, 1835 (1968).
%  %%CITATION = PRLTA,21,1835;%%
%  %14 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  A.~D.~Brody, R.~J.~Cashmore, A.~Kernan, D.~W.~G.~S.~Leith, B.~G.~Levi, B.~C.~Shen, J.~P.~Berge and D.~Herndon {\it et al.},
%  %``Pi- P Elastic Scattering In The Cms Energy Range 1400-mev To 2000-mev,''
%  Phys.\ Rev.\ D {\bf 3}, 2619 (1971).
%  %%CITATION = PHRVA,D3,2619;%%
%  %14 citations counted in INSPIRE as of 27 Nov 201

\bibitem {PDG}J.~Beringer \textit{et al}. (Particle Data Group),
Phys.\ Rev.\ D\textbf{86}, 010001 (2012) and 2013 partial update for the 2014 edition.



\bibitem{theory}
%\cite{}
%\bibitem{} 
  M.~Gell-Mann and MLevy,
  %``The axial vector current in beta decay,''
  Nuovo Cim.\  {\bf 16}, 705 (1960);
  %%CITATION = NUCIA,16,705;%%
  %1479 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  Y.~Nambu and G.~Jona-Lasinio,
  %``Dynamical Model of Elementary Particles Based on an Analogy with Superconductivity. 1.,''
  Phys.\ Rev.\  {\bf 122}, 345 (1961);
  %%CITATION = PHRVA,122,345;%%
  %3845 citations counted in INSPIRE as of 28 Nov 2013
%\cite{}
%\bibitem{} 
  T.~H.~R.~Skyrme,
  %``A Nonlinear field theory,''
  Proc.\ Roy.\ Soc.\ Lond.\ A {\bf 260}, 127 (1961);
  %%CITATION = PRSLA,A260,127;%%
  %1618 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  M.~Gell-Mann,
  %``A Schematic Model of Baryons and Mesons,''
  Phys.\ Lett.\  {\bf 8}, 214 (1964);
  %%CITATION = PHLTA,8,214;%%
  %1752 citations counted in INSPIRE as of 28 Nov 2013
%\cite{}
%\bibitem{} 
  F.~Englert and R.~Brout,
  %``Broken Symmetry and the Mass of Gauge Vector Mesons,''
  Phys.\ Rev.\ Lett.\  {\bf 13}, 321 (1964);
  %%CITATION = PRLTA,13,321;%%
  %2245 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  P.~W.~Higgs,
  %``Broken symmetries, massless particles and gauge fields,''
  Phys.\ Lett.\  {\bf 12}, 132 (1964);
  %%CITATION = PHLTA,12,132;%%
  %2589 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  G.~S.~Guralnik, C.~R.~Hagen and T.~W.~B.~Kibble,
  %``Global Conservation Laws and Massless Particles,''
  Phys.\ Rev.\ Lett.\  {\bf 13}, 585 (1964);
  %%CITATION = PRLTA,13,585;%%
  %1784 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  S.~Weinberg,
  %``Pion scattering lengths,''
  Phys.\ Rev.\ Lett.\  {\bf 17}, 616 (1966);
  %%CITATION = PRLTA,17,616;%%
  %1309 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  L.~D.~Faddeev and V.~N.~Popov,
%  %``Feynman Diagrams for the Yang-Mills Field,''
%  Phys.\ Lett.\ B {\bf 25}, 29 (1967);
% %%CITATION = PHLTA,B25,29;%%
% %1435 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  S.~Weinberg,
%  %``A Model of Leptons,''
%  Phys.\ Rev.\ Lett.\  {\bf 19}, 1264 (1967);
%  %%CITATION = PRLTA,19,1264;%%
%  %8658 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  M.~Gell-Mann, R.~J.~Oakes and B.~Renner,
  %``Behavior of current divergences under SU(3) x SU(3),''
  Phys.\ Rev.\  {\bf 175}, 2195 (1968);
  %%CITATION = PHRVA,175,2195;%%
  %1473 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  S.~R.~Coleman, J.~Wess and B.~Zumino,
  %``Structure of phenomenological Lagrangians. 1.,''
  Phys.\ Rev.\  {\bf 177}, 2239 (1969);
  %%CITATION = PHRVA,177,2239;%%
  %1346 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  J.~S.~Bell and R.~Jackiw,
  %``A PCAC puzzle: pi0 --> gamma gamma in the sigma model,''
  Nuovo Cim.\ A {\bf 60}, 47 (1969);
  %%CITATION = NUCIA,A60,47;%%
  %2343 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  R.~P.~Feynman,
  %``Very high-energy collisions of hadrons,''
  Phys.\ Rev.\ Lett.\  {\bf 23}, 1415 (1969);
  %%CITATION = PRLTA,23,1415;%%
  %1570 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{}
%%\bibitem{} 
%  K.~G.~Wilson,
%  %``Nonlagrangian models of current algebra,''
%  Phys.\ Rev.\  {\bf 179}, 1499 (1969);
%  %%CITATION = PHRVA,179,1499;%%
%  %1639 citations counted in INSPIRE as of 27 Nov 2013
%\cite{Brodsky:1973kr}
%\bibitem{Brodsky:1973kr} 
  S.~J.~Brodsky and G.~R.~Farrar,
  %``Scaling Laws at Large Transverse Momentum,''
  Phys.\ Rev.\ Lett.\  {\bf 31}, 1153 (1973);
  %%CITATION = PRLTA,31,1153;%%
  %1507 citations counted in INSPIRE as of 27 Nov 2013
%\cite{'tHooft:1973jz}
%\bibitem{'tHooft:1973jz} 
  G.~'t Hooft,
  %``A Planar Diagram Theory for Strong Interactions,''
  Nucl.\ Phys.\ B {\bf 72}, 461 (1974);
  %%CITATION = NUPHA,B72,461;%%
  %3517 citations counted in INSPIRE as of 27 Nov 2013
%%\cite{Chodos:1974je}
%%\bibitem{Chodos:1974je} 
%  A.~Chodos, R.~L.~Jaffe, K.~Johnson, C.~B.~Thorn and V.~F.~Weisskopf,
%  %``A New Extended Model of Hadrons,''
%  Phys.\ Rev.\ D {\bf 9}, 3471 (1974);
%  %%CITATION = PHRVA,D9,3471;%%
%  %2148 citations counted in INSPIRE as of 27 Nov 2013
%\cite{Wilson:1974sk}
%\bibitem{Wilson:1974sk} 
  K.~G.~Wilson,
  %``Confinement of Quarks,''
  Phys.\ Rev.\ D {\bf 10}, 2445 (1974);
  %%CITATION = PHRVA,D10,2445;%%
  %3690 citations counted in INSPIRE as of 27 Nov 2013
%\cite{}
%\bibitem{} 
  S.~Weinberg,
  %``Implications of Dynamical Symmetry Breaking,''
  Phys.\ Rev.\ D {\bf 13}, 974 (1976);
  %%CITATION = PHRVA,D13,974;%%
  %1215 citations counted in INSPIRE as of 27 Nov 2013
%\cite{Field:1976ve}
%\bibitem{Field:1976ve} 
  R.~D.~Field and R.~P.~Feynman,
  %``Quark Elastic Scattering as a Source of High Transverse Momentum Mesons,''
  Phys.\ Rev.\ D {\bf 15}, 2590 (1977).
  %%CITATION = PHRVA,D15,2590;%%
  %1083 citations counted in INSPIRE as of 27 Nov 2013




%\cite{Fritzsch:1975sr}
\bibitem{Fritzsch:1975sr} 
  H.~Fritzsch, M.~Gell-Mann and P.~Minkowski,
  %``Vector - Like Weak Currents and New Elementary Fermions,''
  Phys.\ Lett.\ B {\bf 59}, 256 (1975).
  %%CITATION = PHLTA,B59,256;%%
  %318 citations counted in INSPIRE as of 27 Nov 2013


\bibitem{AF}
D.~J.~Gross and F.~Wilczek,
%``Ultraviolet Behavior of Nonabelian Gauge Theories,''
Phys.\ Rev.\ Lett.\ \textbf{30}, 1343 (1973);
%%CITATION = PRLTA,30,1343;%%
%\cite{FERMILAB-PUB-73-049-T}
%\bibitem{FERMILAB-PUB-73-049-T}
D.~J.~Gross and F.~Wilczek,
%``Asymptotically Free Gauge Theories. 1,''
Phys.\ Rev.\ D\ \textbf{8}, 3633 (1973);
%%CITATION = PHRVA,D8,3633;%%
%\cite{94347}
%\bibitem{94347}
H.~D.~Politzer,
%``Setting the scale for predictions of asymptotic freedom,''
Phys.\ Rev.\ D\ \textbf{9}, 2174 (1974);
%%CITATION = PHRVA,D9,2174;%%
%\cite{PRINT-74-1223 (HARVARD)}
%\bibitem{PRINT-74-1223 (HARVARD)}
H.~D.~Politzer,
%``Asymptotic Freedom: An Approach to Strong Interactions,''
Phys.\ Rept.\ \textbf{14}, 129 (1974).
%%CITATION = PRPLC,14,129;%%

%\cite{Parganlija:2012xj}
\bibitem{PhD} 
  D.~Parganlija,
  %``Quarkonium Phenomenology in Vacuum,''
  arXiv:1208.0204 [hep-ph].
  %%CITATION = ARXIV:1208.0204;%%
  %4 citations counted in INSPIRE as of 28 Nov 2013

\bibitem{Jaffeq2q2}R.~L.~Jaffe,
%``Multi-Quark Hadrons. 1. The Phenomenology Of (2 Quark 2 Anti-Quark)
%Mesons,''
Phys.\ Rev.\ D \textbf{15}, 267 (1977).
%%CITATION = PHRVA,D15,267;%%


%\cite{Barnes:1982tx}
\bibitem{Barnes:1982tx} 
  T.~Barnes, F.~E.~Close, F.~de Viron and J.~Weyers,
  %``Q anti-Q G Hermaphrodite Mesons in the MIT Bag Model,''
  Nucl.\ Phys.\ B {\bf 224}, 241 (1983).
  %%CITATION = NUPHA,B224,241;%%
  %210 citations counted in INSPIRE as of 28 Nov 2013

\bibitem{glueballs}
%\cite{Jaffe:1975fd}
%\bibitem{Jaffe:1975fd}
R.~L.~Jaffe and K.~Johnson,
%``Unconventional States Of Confined Quarks And Gluons,''
Phys.\ Lett.\ B \textbf{60}, 201 (1976);
%%CITATION = PHLTA,B60,201;%%
%\cite{Rosenzweig:1981cu}
%\bibitem{Rosenzweig:1981cu}
C.~Rosenzweig, A.~Salomone and J.~Schechter,
%``A Pseudoscalar Glueball, The Axial Anomaly And The Mixing Problem For
%Pseudoscalar Mesons,''
Phys.\ Rev.\ D \textbf{24}, 2545 (1981);
%%CITATION = PHRVA,D24,2545;%%
A.~Salomone, J.~Schechter and T.~Tudron,
%``Properties Of Scalar Gluonium,''
Phys.\ Rev.\ D \textbf{23}, 1143 (1981);
%%CITATION = PHRVA,D23,1143;%%
%\cite{Konoplich:1981ed}
%\bibitem{Konoplich:1981ed}
R.~Konoplich and M.~Shchepkin,
%``Glueballs' Masses In The Bag Model,''
Nuovo Cim.\ A \textbf{67}, 211 (1982);
%%CITATION = NUCIA,67A,211;%%
%\cite{Rosenzweig:1982cb}
%\bibitem{Rosenzweig:1982cb}
C.~Rosenzweig, A.~Salomone and J.~Schechter,
%``How Does A Pseudoscalar Glueball Come Unglued?,''
Nucl.\ Phys.\ B \textbf{206}, 12 (1982) [Erratum-ibid.\ B \textbf{207}, 546
(1982)];
%\cite{Migdal:1982jp}
%\bibitem{Migdal:1982jp}
A.~A.~Migdal and M.~A.~Shifman,
%``Dilaton Effective Lagrangian In Gluodynamics,''
Phys.\ Lett.\ B \textbf{114}, 445 (1982);
%\cite{Jezabek:1982ic}
%%\bibitem{Jezabek:1982ic}
%M.~Jezabek and J.~Szwed,
%%``Glueballs In The Bag Models,''
%Acta Phys.\ Polon.\ B \textbf{14}, 599 (1983);
%%%CITATION = APPOA,B14,599;%%
%\cite{Gomm:1984zq}
%\bibitem{Gomm:1984zq}
H.~Gomm and J.~Schechter,
%``Goldstone Bosons And Scalar Gluonium,''
Phys.\ Lett.\ B \textbf{158}, 449 (1985);
%\cite{Jaffe:1985qp}
%\bibitem{Jaffe:1985qp}
R.~L.~Jaffe, K.~Johnson and Z.~Ryzak,
%``Qualitative Features Of The Glueball Spectrum,''
Annals Phys.\ \textbf{168}, 344 (1986).




%\cite{Parganlija:2012nc}
\bibitem{Hersonissos} 
  D.~Parganlija,
  %``Scalar Mesons and FAIR,''
  J.\ Phys.\ Conf.\ Ser.\  {\bf 426}, 012019 (2013)
  [arXiv:1211.4804 [hep-ph]].
  %%CITATION = ARXIV:1211.4804;%%


\bibitem{barp-p}
C.~Amsler \textit{et al.} [Crystal Barrel
Collaboration],
%``Proton - anti-proton annihilation into eta eta pi: Observation of a scalar resonance decaying into eta eta,''
Phys.\ Lett.\ B \textbf{291}, 347 (1992);
C.~Amsler \textit{et al.} [Crystal Barrel Collaboration],
%``Observation of a scalar resonance decaying to pi+ pi- pi0 pi0 in anti-p p annihilation at rest,''
Phys.\ Lett.\ B \textbf{322}, 431 (1994);
%V.~V.~Anisovich \textit{et al.} [Crystal Barrel Collaboration],
%%``Observation of two J(PC) = 0++ isoscalar resonances at 1365-MeV and 1520-MeV,''
%Phys.\ Lett.\ B \textbf{323}, 233 (1994);
%V.~V.~Anisovich, D.~V.~Bugg, A.~V.~Sarantsev, B.~S.~Zou,
%%``Final state interactions in three meson systems: Analysis of data on anti-p p ---> pi0 pi0 pi0 and eta eta pi0 at rest,''
%Phys.\ Rev.\ D \textbf{50}, 1972 (1994);
D.~V.~Bugg, V.~V.~Anisovich, A.~Sarantsev and B.~S.~Zou,
%``Coupled channel analysis of data on anti-p p $\to$ 3 pi0, eta eta pi0, and
%eta pi0 pi0 at rest, with the N/D method,''
Phys.\ Rev.\ D \textbf{50}, 4412 (1994);
C.~Amsler \textit{et al.} [Crystal Barrel
Collaboration],
%``eta eta-prime threshold enhancement in anti-p p annihilations into pi0 eta eta-prime at rest,''
Phys.\ Lett.\ B \textbf{340}, 259 (1994);
%C.~Amsler, D.~S.~Armstrong, I.~Augustin, C.~A.~Baker,
%B.~M.~Barnett, C.~J.~Batty, K.~Beuchert, P.~Birien \textit{et al.},
%%``High statistics study of f0 (1500) decay into pi0 pi0,''
%Phys.\ Lett.\ B \textbf{342}, 433 (1995);
%C.~Amsler \textit{et al.} [Crystal Barrel
%Collaboration],
%%``High statistics study of f0 (1500) decay into eta eta,''
%Phys.\ Lett.\ B \textbf{353}, 571 (1995);
%\bibitem {Amsler:1995bf}
C.~Amsler \textit{et al.} [Crystal Barrel Collaboration],
%``Coupled channel analysis of anti-p p annihilation into pi0 pi0 pi0, pi0 eta eta and pi0 pi0 eta,''
Phys.\ Lett.\ B \textbf{355}, 425 (1995);
%A.~Abele, J.~Adomeit, D.~S.~Armstrong, C.~A.~Baker,
%C.~J.~Batty, M.~Benayoun, A.~Berdoz, K.~Beuchert \textit{et al.},
%%``Further analysis of anti-p p ---> 3pi0, eta eta pi0 and eta pi0 pi0 at rest,''
%Nucl.\ Phys.\ A \textbf{609}, 562 (1996);
A.~Bertin {\it et al.} [OBELIX Collaboration],
%``Spin-parity analysis of the final state pi+ pi- pi0 from anti-p p annihilation at rest in hydrogen targets at three densities,''
Phys.\ Lett.\ B {\bf 408}, 476 (1997);
%A.~V.~Anisovich \textit{et al.},
%%``Observation of f0(1770) --> eta eta in anti-p p --> eta eta pi0  reactions
%%from 600-MeV/c to 1200-MeV/c,''
%Phys.\ Lett.\ B \textbf{449}, 154 (1999);
C.~Amsler \textit{et al.} [Crystal Barrel Collaboration],
%``Study of K anti-K resonances in anti-p p ---> K+ K- pi0 at 900-MeV/c and 1640-MeV/c,''
Phys.\ Lett.\ B \textbf{639}, 165 (2006).



\bibitem{p-p} T.~Akesson \textit{et al.} [Axial Field Spectrometer
Collaboration],
%``A Search For Glueballs And A Study Of Double Pomeron Exchange At The Cern Intersecting Storage Rings,''
Nucl.\ Phys.\ B\ \textbf{264}, 154 (1986);
%%CITATION = NUPHA,B264,154;%%
%\cite{CERN-EP/86-11}
%%\bibitem{CERN-EP/86-11}
%A.~Breakstone \textit{et al.} [Ames-Bologna-CERN-Dortmund-Heidelberg-Warsaw
%Collaboration],
%%``Production Of The F0 Meson In The Double Pomeron Exchange Reaction P P ---> P P Pi+ Pi- At S**(1/2) = 62-gev,''
%Z.\ Phys.\ C\ \textbf{31}, 185 (1986);
%T.~A.~Armstrong \textit{et al.} [WA76
%Collaboration],
%%``Evidence For New States Produced In The Central Region In The Reaction P P ---> P(f) (pi+ Pi- Pi+ Pi-) P(s) At 300-gev/c,''
%Phys.\ Lett.\ B \textbf{228}, 536 (1989);
%%CITATION = ZEPYA,C31,185;%%
%\cite{Breakstone:1990at}
%\bibitem{Breakstone:1990at}
A.~Breakstone \textit{et al.} [Ames-Bologna-CERN-Dortmund-Heidelberg-Warsaw
Collaboration],
%``The Reaction pomeron-pomeron ---> pi+ pi- and an unusual production mechanism for the f2 (1270),''
Z.\ Phys.\ C \textbf{48}, 569-576 (1990);
%S.~Abatzis \textit{et al.} [WA91
%Collaboration],
%%``Observation of a narrow scalar meson at 1450-MeV in the reaction p p ---> p(f) (pi+ pi- pi+ pi-) p(s) at 450-GeV/c using the CERN Omega Spectrometer,''
%Phys.\ Lett.\ B \textbf{324}, 509 (1994);
%\cite{Antinori:1995wz}
%\bibitem{Antinori:1995wz}
F.~Antinori \textit{et al.} [WA91 Collaboration],
%``A Further study of the centrally produced pi+ pi- and pi+ pi- pi+ pi- channels in p p interactions at 300-GeV/c and 450-GeV/c,''
Phys.\ Lett.\ B \textbf{353}, 589 (1995);
D.~Alde \textit{et al.} [GAMS Collaboration],
%``Study of the pi0 pi0 system in p p central collisions at 450-GeV/c,''
Phys.\ Lett.\ B \textbf{397}, 350 (1997);
%D.~Barberis \textit{et al.} [WA102 Collaboration],
%%``A Study of the centrally produced pi+ pi- pi+ pi- channel in p p interactions at 450-GeV/c,''
%Phys.\ Lett.\ B \textbf{413}, 217 (1997) [hep-ex/9707021];
%D.~Barberis \textit{et al.} [WA102 Collaboration],
%%``A Partial wave analysis of the centrally produced K+ K- and K0(S) K0(S) systems in p p interactions at 450-GeV/c and new information on the spin of the f(J)(1710),''
%Phys.\ Lett.\ B \textbf{453}, 305 (1999);
D.~Barberis \textit{et al.} [WA102 Collaboration],
%``A Partial wave analysis of the centrally produced pi+ pi- system in p p interactions at 450-GeV/c,''
Phys.\ Lett.\ B \textbf{453}, 316 (1999) [arXiv:hep-ex/9903043];
R.~Bellazzini \textit{et al.} [GAMS Collaboration],
%``A partial wave analysis of the centrally produced pi0 pi0 system in p p interactions at 450-GeV/c,''
Phys.\ Lett.\ B \textbf{467}, 296 (1999).

\bibitem{WA76} T.~A.~Armstrong \textit{et al.} [WA76
Collaboration],
%``Study of the centrally produced pi pi and K anti-K systems at 85-GeV/c and 300-GeV/c,''
Z.\ Phys.\ C \textbf{51}, 351 (1991).

\bibitem{e+e-} 
C.~Edwards, R.~Partridge, C.~Peck, F.~Porter,
D.~Antreasyan, Y.~F.~Gu, W.~S.~Kollmann, M.~Richardson \textit{et al.},
%``Observation of an eta eta Resonance in J/psi Radiative Decays,''
Phys.\ Rev.\ Lett.\ \textbf{48}, 458 (1982);
R.~M.~Baltrusaitis \textit{et al.} [MARK-III
Collaboration],
%``A Study Of The Radiative Decay J / Psi $\to$ Gamma Rho Rho,''
Phys.\ Rev.\ D \textbf{33}, 1222 (1986);
%%CITATION = PHRVA,D33,1222;%%
%\cite{Bolton:1991nx}
%\bibitem{Bolton:1991nx}
T.~Bolton \textit{et al.},
%``Observation of f-1(1285) $\to$ pi+ pi- pi+ pi- in radiative J / psi
%decays,''
Phys.\ Lett.\ B \textbf{278}, 495 (1992);
R.~R.~Akhmetshin \textit{et al.} [CMD-2
Collaboration],
%``Study of the phi decays into pi0 pi0 gamma and eta pi0 gamma final states,''
Phys.\ Lett.\ B \textbf{462}, 380 (1999) [arXiv:hep-ex/9907006];
%J.~Z.~Bai \textit{et al.} [BES Collaboration],
%%``Partial Wave Analysis of $J/\psi\to\gamma(\pi^+\pi^-\pi^+\pi^-)$,''
%Phys.\ Lett.\ B \textbf{472}, 207 (2000) [arXiv:hep-ex/9909040];
%B.~S.~Zou,
%%``Light hadron spectroscopy at BEPC,''
%Nucl.\ Phys.\ A \textbf{692}, 362 (2001) [arXiv:hep-ph/0011174];
 M.~Ablikim {\it et al.} [BES Collaboration],
  %``The sigma pole in J / psi ---> omega pi+ pi-,''
  Phys.\ Lett.\ B {\bf 598}, 149 (2004)
  [hep-ex/0406038];
  M.~Ablikim \textit{et al.} [BES Collaboration],
%``Resonances in J / psi ---> phi pi+ pi- and phi K+ K-,''
Phys.\ Lett.\ B \textbf{607}, 243 (2005) [arXiv:hep-ex/0411001];
%M.~Ablikim \textit{et al.} [BES Collaboration],
%%``Partial Wave Analysis of $\chi_{c0}\to\pi^+\pi^-K^+K^-$,''
%Phys.\ Rev.\ D \textbf{72}, 092002 (2005) [arXiv:hep-ex/0508050];
%%CITATION = PHLTA,B607,243;%
F.~Ambrosino \textit{et al.} [KLOE Collaboration],
%``Dalitz plot analysis of e+ e- ---> pi0 pi0 gamma events at s**(1/2) approximately M(phi) with the KLOE detector,''
Eur.\ Phys.\ J.\ C \textbf{49}, 473 (2007) [hep-ex/0609009].



\bibitem{pi-N} 
B.~Hyams, C.~Jones, P.~Weilhammer, W.~Blum, H.~Dietl,
G.~Grayer, W.~Koch, E.~Lorenz \textit{et al.},
%``pi pi Phase Shift Analysis from 600-MeV to 1900-MeV,''
Nucl.\ Phys.\ B \textbf{64}, 134 (1973);
%D.~M.~Binnie, J.~Carr, N.~C.~Debenham, A.~Duane,
%D.~A.~Garbutt, W.~G.~Jones, J.~Keyne, I.~Siotis \textit{et al.},
%%``Direct evidence for the s* meson near the k anti-k threshold,''
%Phys.\ Rev.\ Lett.\ \textbf{31}, 1534 (1973);
H.~Becker \textit{et al.} [CERN-Cracow-Munich
Collaboration],
%``A Model Independent Partial Wave Analysis Of The Pi+ Pi- System Produced At
%Low Four Momentum Transfer In The Reaction Pi- P (Polarized) $\to$ Pi+ Pi- N
%At 17.2-Gev/C,''
Nucl.\ Phys.\ B \textbf{151}, 46 (1979);
D.~Alde \textit{et al.} [Serpukhov-Brussels-Los
Alamos-Annecy(LAPP) Collaboration],
%``OBSERVATION OF A 1750-MeV NARROW MESON DECAYING INTO eta eta,''
Phys.\ Lett.\ B \textbf{182}, 105 (1986);
%D.~Alde \textit{et al.} [IFVE-Brussels-Annecy-Los Alamos Collaboration],
%%``Production mechanism of the X (1740) meson,''
%Phys.\ Lett.\ \textbf{B284}, 457 (1992);
%D.~Alde \textit{et al.} [GAMS Collaboration],
%%``Study of the f0 (995) resonance in the pi0 pi0 decay channel,''
%Z.\ Phys.\ C \textbf{66}, 375 (1995);
D.~Alde \textit{et al.} [GAMS Collaboration],
%``Study of the pi0 pi0 system with the GAMS-4000 spectrometer at 100-GeV/c,''
Eur.\ Phys.\ J.\ A \textbf{3}, 361 (1998);
E.~M.~Aitala {\it et al.} [E791 Collaboration],
  %``Experimental evidence for a light and broad scalar resonance in D+ ---> pi- pi+ pi+ decay,''
  Phys.\ Rev.\ Lett.\ {\bf 86}, 770 (2001)
  [hep-ex/0007028].



\bibitem{models}
%\bibitem{Minkowski:1998mf}
P.~Minkowski and W.~Ochs,
%``Identification of the glueballs and the scalar meson nonet of lowest
%mass,''
Eur.\ Phys.\ J.\ C \textbf{9}, 283 (1999) [arXiv:hep-ph/9811518];
F.~Giacosa, T.~.Gutsche, V.~E.~Lyubovitskij and A.~Faessler,
%``Scalar meson and glueball decays within a effective chiral approach,''
  Phys.\ Lett.\ B {\bf 622}, 277 (2005)
  [hep-ph/0504033];
%  F.~Giacosa, T.~.Gutsche, V.~E.~Lyubovitskij and A.~Faessler,
%  ``Scalar nonet quarkonia and the scalar glueball: Mixing and decays in an
%  effective chiral approach,''
  Phys.\ Rev.\ D {\bf 72}, 094006 (2005)
  [hep-ph/0509247];
%  F.~Giacosa, T.~.Gutsche, V.~E.~Lyubovitskij and A.~Faessler,
%  ``Decays of tensor mesons and the tensor glueball in an effective field
% approach,''
  Phys.\ Rev.\ D {\bf 72}, 114021 (2005)
  [hep-ph/0511171];
%\cite{Parganlija:2010fz}
%\bibitem{Parganlija:2010fz} 
  D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
  %``Vacuum Properties of Mesons in a Linear Sigma Model with Vector Mesons and Global Chiral Invariance,''
  Phys.\ Rev.\ D {\bf 82}, 054024 (2010)
  [arXiv:1003.4934 [hep-ph]];
  %%CITATION = ARXIV:1003.4934;%%
  %46 citations counted in INSPIRE as of 02 Dec 2013
%\cite{Janowski:2011gt}
%\bibitem{Janowski:2011gt} 
  S.~Janowski, D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
  %``The Glueball in a Chiral Linear Sigma Model with Vector Mesons,''
  Phys.\ Rev.\ D {\bf 84}, 054007 (2011)
  [arXiv:1103.3238 [hep-ph]];
  %%CITATION = ARXIV:1103.3238;%%
  %22 citations counted in INSPIRE as of 02 Dec 2013
%\cite{Chatzis:2011qz}
%\bibitem{Chatzis:2011qz} 
  P.~Chatzis, A.~Faessler, T.~Gutsche and V.~E.~Lyubovitskij,
  %``Hadronic and radiative three-body decays of J/psi involving the scalars f0(1370), f0(1500) and f0(1710),''
  Phys.\ Rev.\ D {\bf 84}, 034027 (2011)
  [arXiv:1105.1676 [hep-ph]];
  %%CITATION = ARXIV:1105.1676;%%
  %1 citations counted in INSPIRE as of 19 Dec 2013
%\cite{Parganlija:2012fy}
%\bibitem{Parganlija:2012fy} 
  D.~Parganlija, P.~Kovacs, G.~Wolf, F.~Giacosa and D.~H.~Rischke,
  %``Meson vacuum phenomenology in a three-flavor linear sigma model with (axial-)vector mesons,''
  Phys.\ Rev.\ D {\bf 87}, 014011 (2013)
  [arXiv:1208.0585 [hep-ph]].
  %%CITATION = ARXIV:1208.0585;%%
  %27 citations counted in INSPIRE as of 02 Dec 2013

\bibitem{lattice}
C.~Morningstar and M.~J.~Peardon,
%``Simulating the scalar glueball on the lattice,''
AIP Conf.\ Proc.\ \textbf{688}, 220 (2004) [arXiv:nucl-th/0309068];
%%CITATION = APCPC,688,220;%%
%\cite{Loan:2005ff}
%\bibitem{Loan:2005ff}
M.~Loan, X.~Q.~Luo and Z.~H.~Luo,
%``Monte Carlo study of glueball masses in the Hamiltonian limit of SU(3)
%lattice gauge theory,''
Int.\ J.\ Mod.\ Phys.\ A \textbf{21}, 2905 (2006) [arXiv:hep-lat/0503038];
%%CITATION = IMPAE,A21,2905;%%
%\cite{Gregory:2005yr}
%\bibitem{Gregory:2005yr}
E.~B.~Gregory, A.~C.~Irving, C.~C.~McNeile, S.~Miller and Z.~Sroczynski,
%``Scalar glueball and meson spectroscopy in unquenched lattice QCD with
%improved staggered quarks,''
PoS \textbf{LAT2005}, 027 (2006) [arXiv:hep-lat/0510066];
%\cite{Chen:2005mg}
%\bibitem{Chen:2005mg}
Y.~Chen \textit{et al.},
%``Glueball spectrum and matrix elements on anisotropic lattices,''
Phys.\ Rev.\ D \textbf{73}, 014516 (2006) [arXiv:hep-lat/0510074].
%%CITATION = PHRVA,D73,014516;%%


\bibitem{holography}
%\cite{Csaki:1998qr}
%\bibitem{Csaki:1998qr} 
  C.~Csaki, H.~Ooguri, Y.~Oz and J.~Terning,
  %``Glueball mass spectrum from supergravity,''
  JHEP {\bf 9901}, 017 (1999)
  [hep-th/9806021];
  %%CITATION = HEP-TH/9806021;%%
  %253 citations counted in INSPIRE as of 22 May 2013
%  %\cite{de Mello Koch:1998qs}
%%\bibitem{de Mello Koch:1998qs} 
%  R.~de Mello Koch, A.~Jevicki, M.~Mihailescu and J.~P.~Nunes,
%  %``Evaluation of glueball masses from supergravity,''
%  Phys.\ Rev.\ D {\bf 58}, 105009 (1998)
%  [hep-th/9806125];
  %%CITATION = HEP-TH/9806125;%%
  %127 citations counted in INSPIRE as of 22 May 2013
  %\cite{Csaki:1998cb}
%%\bibitem{Csaki:1998cb} 
%  C.~Csaki, Y.~Oz, J.~Russo and J.~Terning,
%  %``Large N QCD from rotating branes,''
%  Phys.\ Rev.\ D {\bf 59}, 065012 (1999)
%  [hep-th/9810186];
%  %%CITATION = HEP-TH/9810186;%%
% %73 citations counted in INSPIRE as of 23 May 2013
R.~C.~Brower, S.~D.~Mathur and C.~-I Tan,
  %``Glueball spectrum for QCD from AdS supergravity duality,''
  Nucl.\ Phys.\ B {\bf 587}, 249 (2000)
  [hep-th/0003115];
  %\cite{BoschiFilho:2002ta}
%%\bibitem{BoschiFilho:2002ta} 
%  H.~Boschi-Filho and N.~R.~F.~Braga,
%  %``QCD / string holographic mapping and glueball mass spectrum,''
%  Eur.\ Phys.\ J.\ C {\bf 32}, 529 (2004)
%  [hep-th/0209080];
%JHEP {\bf 0305}, 009 (2003)
% [hep-th/0212207];
% %%CITATION = HEP-TH/0209080;%%
% %88 citations counted in INSPIRE as of 22 May 2013
%  %\cite{BoschiFilho:2002vd}
%%\bibitem{BoschiFilho:2002vd} 
%  H.~Boschi-Filho and N.~R.~F.~Braga,
%  %``Gauge / string duality and scalar glueball mass ratios,''
%  JHEP {\bf 0305}, 009 (2003)
%  [hep-th/0212207];
%  %%CITATION = HEP-TH/0212207;%%
%  %113 citations counted in INSPIRE as of 22 May 2013
  %\cite{Ametller:2003dj}
%%\bibitem{Ametller:2003dj} 
%  L.~Ametller, J.~M.~Pons and P.~Talavera,
%  %``On the consistency of the N=1 SYM spectra from wrapped five-branes,''
%  Nucl.\ Phys.\ B {\bf 674}, 231 (2003)
%  [hep-th/0305075];
%  %%CITATION = HEP-TH/0305075;%%
  %16 citations counted in INSPIRE as of 22 May 2013
  %\cite{Apreda:2003sy}
%%\bibitem{Apreda:2003sy} 
%  R.~Apreda, D.~E.~Crooks, N.~J.~Evans and M.~Petrini,
%  %``Confinement, glueballs and strings from deformed AdS,''
%  JHEP {\bf 0405}, 065 (2004)
% [hep-th/0308006];
%  %%CITATION = HEP-TH/0308006;%%
%  %26 citations counted in INSPIRE as of 23 May 2013
% %\cite{deTeramond:2005su}
%\bibitem{deTeramond:2005su} 
  G.~F.~de Teramond and S.~J.~Brodsky,
  %``Hadronic spectrum of a holographic dual of QCD,''
  Phys.\ Rev.\ Lett.\  {\bf 94}, 201601 (2005)
  [hep-th/0501022];
  %%CITATION = HEP-TH/0501022;%%
  %250 citations counted in INSPIRE as of 22 May 2013
  %\cite{Evans:2005ip}
%\bibitem{Evans:2005ip} 
  N.~Evans, J.~P.~Shock and T.~Waterson,
  %``Towards a perfect QCD gravity dual,''
  Phys.\ Lett.\ B {\bf 622}, 165 (2005)
  [hep-th/0505250];
  %%CITATION = HEP-TH/0505250;%%
  %28 citations counted in INSPIRE as of 22 May 2013
  %\cite{BoschiFilho:2005yh}
%\bibitem{BoschiFilho:2005yh} 
  H.~Boschi-Filho, N.~R.~F.~Braga and H.~L.~Carrion,
  %``Glueball Regge trajectories from gauge/string duality and the Pomeron,''
  Phys.\ Rev.\ D {\bf 73}, 047901 (2006)
  [hep-th/0507063].
  %%CITATION = HEP-TH/0507063;%%
  %71 citations counted in INSPIRE as of 23 May 2013
  %\cite{Colangelo:2007pt}
%%\bibitem{Colangelo:2007pt} 
%  P.~Colangelo, F.~De Fazio, F.~Jugeau and S.~Nicotri,
%  %``On the light glueball spectrum in a holographic description of QCD,''
%  Phys.\ Lett.\ B {\bf 652}, 73 (2007)
%  [hep-ph/0703316];
%  %%CITATION = HEP-PH/0703316;%%
%  %72 citations counted in INSPIRE as of 23 May 2013
  %\cite{Gursoy:2007er}
%\bibitem{Gursoy:2007er} 
 % U.~Gursoy, E.~Kiritsis and F.~Nitti,
%  %``Exploring improved holographic theories for QCD: Part II,''
%  JHEP {\bf 0802}, 019 (2008)
%  [arXiv:0707.1349 [hep-th]];
%  %%CITATION = ARXIV:0707.1349;%%
%  %165 citations counted in INSPIRE as of 22 May 2013
% %\cite{Forkel:2007ru}
%%\bibitem{Forkel:2007ru} 
%  H.~Forkel,
%  %``Holographic glueball structure,''
%  Phys.\ Rev.\ D {\bf 78}, 025001 (2008)
% [arXiv:0711.1179 [hep-ph]]



%\cite{Barate:1999ze}
\bibitem{ALEPH} 
  R.~Barate {\it et al.}  [ALEPH Collaboration],
  %``Search for the glueball candidates f(0)(1500) and f(J)(1710) in gamma gamma collisions,''
  Phys.\ Lett.\ B {\bf 472}, 189 (2000)
  [hep-ex/9911022].
  %%CITATION = HEP-EX/9911022;%%
  %43 citations counted in INSPIRE as of 02 Dec 2013

%\cite{Acciarri:2000ex}
\bibitem{L3} 
  M.~Acciarri {\it et al.}  [L3 Collaboration],
  %``$K^0_{S} K^0_{S}$ final state in two photon collisions and implications for glueballs,''
  Phys.\ Lett.\ B {\bf 501}, 173 (2001)
  [hep-ex/0011037].
  %%CITATION = HEP-EX/0011037;%%
  %49 citations counted in INSPIRE as of 02 Dec 2013

%\bibitem{BESII} 
\bibitem {BESII2004}
M.~Ablikim \textit{et al.} [BES Collaboration],
%``Resonances in J / psi ---> phi pi+ pi- and phi K+ K-,''
Phys.\ Lett.\ B \textbf{607}, 243 (2005) [arXiv:hep-ex/0411001].
%%CITATION = PHLTA,B607,243;%%

%\cite{Lutz:2009ff}
\bibitem{PANDA} 
  M.~F.~M.~Lutz {\it et al.}  [PANDA Collaboration],
  %``Physics Performance Report for PANDA: Strong Interaction Studies with Antiprotons,''
  arXiv:0903.3905 [hep-ex].
  %%CITATION = ARXIV:0903.3905;%%
  %166 citations counted in INSPIRE as of 02 Dec 2013



\end{thebibliography}

\documentclass[a4paper]{jpconf}
\usepackage{graphicx}
\begin{document}
\title{Mesons, PANDA and the scalar glueball}

\author{Denis Parganlija}

\address{Vienna University of Technology, Institute for Theoretical Physics\\Wiedner Hauptstr.\ 8-10/E136, 1040 Vienna, Austria}

\ead{denisp@hep.itp.tuwien.ac.at}

\begin{abstract}
The non-perturbative nature of Quantum Chromodynamics (QCD) at low energies has prompted the expectation that the gauge-bosons of QCD -- gluons --
might give rise to compound objects denoted as glueballs. Experimental signals for glueballs have represented a matter of research for various
collaborations in the last decades; future research in this direction is a main endeavour planned by the PANDA Collaboration at FAIR. Hence in this article I review
some of the outstanding issues in the glueball search, particularly with regard to the ground state -- the scalar glueball, and discuss the relevance for PANDA at FAIR.
\end{abstract}

\section{Introduction}
Approximately five decades ago a truly remarkable number of new, strongly interacting resonances was discovered leading way to a new direction in the development
of nuclear physics (a non-exhaustive list of the relevant articles is presented in Ref.\ \cite{6d}
and further work can be traced in the Particle Data Tables \cite{PDG}). These discoveries
coincided with the appearance of many renowned articles in theoretical physics -- a selection of which can be found in Ref.\ \cite{theory} --
that represented a basis for the development of a theory of strong interactions. The most viable approach was subsequently found to be the one where
``the strong interactions are described by an unbroken gauge theory based on
tricolored quarks and color octet vector gluons, with color supposed to be entirely
confined, so that quarks are fractionally charged'' with the suggestion that ``a good name for this theory
is quantum chromodynamics (QCD)'' \cite{Fritzsch:1975sr}.\\
\\
Quantum Chromodynamics has various interesting features, one of the most important of which is a scale-dependent coupling \cite{AF} of such nature that, at higher energies,
it leads to weaker interactions between the relevant degrees of freedom -- quarks and gluons. Conversely, at lower energies, quarks and gluons are confined into
more macroscopic structures: hadrons. According to the total spin $J$, hadrons are divided into mesons (that possess integer spin) and baryons (that possess half-integer spin). Determination of
hadron properties from Quantum Chromodynamics is a highly non-trivial task due to the inapplicability of perturbative approaches in the energy region where hadrons emerge.
A circumvention of this problem is introduced by effective models containing the so-called ``constituent'' quarks and gluons that, in the particular case of mesons, allow for the particle 
classification as quarkonia ($\bar{q}q$, see Ref.\ \cite{PhD} and references therein), tetraquarks ($\bar{q}\bar{q} qq$ \cite{Jaffeq2q2}), hybrids 
(constituent quarks with excited gluon degrees of freedom \cite{Barnes:1982tx}) and others.
\\
%\\
The non-Abelian nature of QCD implies gluon self-interaction that is governed by the same scale-dependent coupling as in the
quark sector. Hence the natural expectation is that, analogously to quarks, the gluons also bind into composite objects -- the glueballs \cite{glueballs}.
\\
\\
There are at least two important reasons to consider glueballs as low-energy degrees of freedom in QCD:
\begin{itemize}
\item Glueball mass is generated exclusively by the gluon self-interaction. The Higgs
mechanism does not contribute to the mass generation of glueballs rendering them uniquely suitable
for the research of the strong interaction.

\item Gluons are vector particles and hence glueballs possess integer spin. This implies that glueballs belong
to the meson spectrum -- and therefore the meson spectrum can only be complete if the spectrum of glueballs is known.

\end{itemize}
As already noted, glueballs have long been subject to experimental and theoretical research.
%\section{Glueballs, experiment and theory}
In general, low-energy mesons are mainly produced in these channels \cite{PhD,Hersonissos}:
\begin{itemize}

\item antiproton-proton ($\bar{p}$-$p$) collisions \cite{barp-p}

\item proton-proton ($p$-$p$) collisions \cite{p-p,WA76}

\item electron-positron annihilation into $J/\psi$ or $\varphi(1020)$ and their subsequent decay into further meson
resonances \cite{e+e-}

\item pion-nucleon scattering reactions \cite{WA76,pi-N}.

\end{itemize}
A glueball state is expected to be preferably produced in radiative decays and to possess a strongly suppressed $\gamma \gamma$
decay (or a strongly suppressed production in the $\gamma \gamma$ collisions).\\
\\
On the theory side, there are various means suitable for the glueball identification. These range from
effective models \cite{glueballs,models} to
lattice QCD \cite{lattice} and holographic-QCD approaches \cite{holography}.
\\
\\
However, a clear experimental signal for a glueball can only be obtained if the resonance has a strongly suppressed mixing to other states with 
the same quantum numbers.
%This implies, for example, that the so-called exotic mesons -- i.e., those with quantum numbers inaccessible
%to $\bar{q}q$ configurations -- suffer only marginally from mixing effects rendering relatively clear signals.
Otherwise the glueball search can become massively complicated, as we illustrate in the following on the example of the scalar glueball.

\section{The scalar glueball}

Scalar mesons are QCD degrees of freedom that possess quantum numbers $J^{PC} = 0^{++}$,
where $J$ is the total spin, $P$ represents parity and $C$ is the charge 
conjugation; mesons with isospin $I=0$ are denoted as isoscalars.
\\
PDG data suggest the existence of 
five $IJ^{PC} = 00^{++}$ states in the energy region up to 1.8 GeV: $f_0(500)$ or $\sigma$, $f_0(980)$, $f_0(1370)$, $f_0(1500)$ and $f_0(1710)$ \cite{PDG}.
At the level of quantum numbers, all the mentioned states are viable candidates for the scalar glueball. %The expectation is that
Then the glueball
should be identified by analysing ({\it i}) the production mechanism and ({\it ii}) the decay channels.\\
\\
For the production mechanism, the situation is as follows \cite{PDG}:
\begin{itemize}
 \item Produced in radiative decays: fulfilled by $f_0(980)$, $f_0(1500)$ and $f_0(1710)$.
 
 \item Not seen in $\gamma \gamma$ collisions: fulfilled by $f_0(1500)$ and $f_0(1710)$.
 
\end{itemize}
Then $f_0(1500)$ and $f_0(1710)$ appear to be the strongest candidates for the scalar glueball since they fulfill both of the above criteria.
At this point, let us briefly comment on the above $\gamma \gamma$ data since they are of great importance for the glueball search and also
since they allow us to illustrate experimental ambiguities in the $f_0$ sector. The main reason to include
$f_0(1500)$ and $f_0(1710)$ into the above list is the fact that there was a notable absence of both of these states in the $\gamma\gamma$ data analysed by the ALEPH
Collaboration in 2000 \cite{ALEPH}.
Contrarily, an independent search by the L3 Collaboration -- that approximately coincided with the ALEPH results -- observed a clear signal
for $f_0(1710)$ while no signal was found for $f_0(1500)$ \cite{L3}. These are currently the best, and the latest, experimental data containing
negative results on photon coupling of the $f_0$ states, but they do not appear to be conclusive.
\\
\\
Hence the expectation might be that the decay channels of the $f_0$ resonances should provide us with more clear-cut results for the glueball identification.
The reason is that, in the case of the unbroken flavour symmetry, one can calculate exact decay-width ratios valid for a glueball state without mixing
to states containing quarks (``pure  glueball''). Then, for example, the ratio of the pion decay width to the kaon decay width is 3/4.\\
Unfortunately, neither the data on $f_0(1500)$ nor those on $f_0(1710)$ follow this expectation:
\begin{itemize}
 \item The $f_0(1500)$ resonance decays predominantly into pions ($2\pi$ and $4\pi$) whereas the kaon decay is suppressed;
the ratio $\Gamma_{f_0(1500)\rightarrow \pi\pi} / \Gamma_{f_0(1500)\rightarrow KK} \simeq 4$ \cite{PDG}.

\item The $f_0(1710)$ resonance decays predominantly into kaons and the pion decay is suppressed; the ratio
$\Gamma_{f_0(1710)\rightarrow \pi\pi} / \Gamma_{f_0(1710)\rightarrow KK} \simeq (0.2 - 0.4)$, depending on collaboration \cite{PhD}.

\end{itemize}
These results demonstrate clearly that there is no exclusive glueball state in the scalar sector 
-- one can only identify a resonance
that is \emph{predominantly} a glueball. Then, as already indicated, various model calculations have to be utilised in order to study mixing patterns of the pure glueball
with $\bar{q}q$, $\bar{q}\bar{q}qq$ and other states \cite{glueballs,models}. Lattice results \cite{lattice} are also of extreme importance, especially since the inclusion
of dynamical quarks has become more prominent; realistic approaches in holographic QCD can be very helpful to the cause as well \cite{holography}.
\\
\\
However, there is a further complication: the possible existence of a new scalar state, $f_0(1790)$.

\section{The \boldmath $f_0(1790)$ resonance and PANDA}

The BES II Collaboration claimed the existence of the new $f_0(1790)$ resonance in 2004 \cite{BESII2004}. The mass and the decay width of the 
resonance were determined by the Collaboration to be $m_{f_{0}(1790)}=1790_{-30}^{+40}$ and $\Gamma_{f_{0}(1790)}=270_{-30}^{+60}$ MeV, respectively.
The stated large decay width implies a strong overlap
with $f_0(1710)$. However, there is a clear point of distinction between the two resonances: as already mentioned, $f_0(1710)$ is reconstructed predominantly in the kaon
decay channels whereas $f_0(1790)$ is reconstructed predominantly in the pion decay channels.\\\\
The distinction between $f_0(1710)$ and $f_0(1790)$ is further ascertained by their production mechanism that involves $J/\psi$ decays \cite{BESII2004}:
({\it i}) $J/\psi\rightarrow\varphi K^+ K^-$;
({\it ii}) $J/\psi\rightarrow\varphi \pi^{+}\pi^{-}$;
({\it iii}) $J/\psi\rightarrow\omega K^{+}K^{-}$;
({\it iv}) $J/\psi\rightarrow\omega \pi^{+}\pi^{-}$.
The $f_0(1710)$ resonance is reconstructed from the decays ({\it i}) and ({\it iii})
whereas the decays ({\it ii}) and ({\it iv}) allow for the reconstruction of $f_{0}(1790)$.
Then assuming $f_{0}(1710)$ and $f_{0}(1790)$ to be the same resonance leads to a contradiction: 
such a resonance would have to possess a pion-to-kaon-decay ratio of $1.82 \pm 0.33$ according to reactions
({\it i}) and ({\it ii}) and, simultaneously, the pion-to-kaon-decay ratio $< 0.11$ according to reactions
({\it iii}) and ({\it iv}) \cite{BESII2004}. That is obviously not possible -- and hence
the BES II data suggest $f_{0}(1710)$ and $f_{0}(1790)$ to represent two distinct resonances.
\\
\\
Let us now discuss the importance of this statements for the PANDA experiments at FAIR. The PANDA Collaboration intends to
study interactions between antiprotons and fixed-target protons and nuclei in the momentum range of 1.5-15 GeV
with the stated goal of
exploring the spectroscopy of glueballs, multiquark and exotic states \cite{PANDA}.
\\\\
The search for glueballs is an extremely important endeavour of the Collaboration since, as already indicated, these states offer a unique 
insight into the strong interaction. However, there are numerous complications along this path, exhibited for example in the case of the ground state
-- the scalar glueball.
\\
This state will inevitably have strong mixing with mesons of $\bar{q}q$, $\bar{q}\bar{q}qq$ and other structures.
Nonetheless, it is extremely important to identify the scalar glueball since otherwise the glueball spectrum would not be complete.
In this case, however, one must bear in mind that
the most basic condition for the identification is to ascertain
\emph{the number} of the scalar states.
In the light of the BES II data, it is then of extreme importance for the PANDA Collaboration
to search for, and confirm -- or disprove -- the existence of, the $f_0(1790)$ resonance since, if this resonance exists, it will inevitably overlap
with the scalar glueball.

\section{Conclusions}
Glueball search is one of the most interesting challenges for the future PANDA experiments. These particles are expected to
be generated by the strong interaction and therefore the determination of their spectrum is of extreme importance for our understanding
of low-energy QCD. The determination of the ground, scalar, state is imperative for the completeness of the glueball spectrum.
Due to the mixing of the pure glueball state with other mesons that have the same quantum numbers, the scalar glueball can only be 
reliably identified if all scalar mesons ($f_0$ resonances) are known. At present, it is not clear whether this is the case: the existence of a new
resonance denoted as $f_0(1790)$ has been suggested by the BES II data
locating $f_0(1790)$ exactly in the energy region where the scalar glueball is expected.
Hence it is very important for the existence of
$f_0(1790)$ to be confirmed or disproved -- and the 
PANDA Collaboration is in a unique position to perform that task.

\ack
I am very grateful to D.\ Bugg, F.\ Giacosa, S.\ Janowski, P.\ Kovacs, A.\ Rebhan, D.\ H.\ Rischke and Gy.\ Wolf for extensive discussions
regarding meson spectroscopy.

\end{document}